1 Find thevalues of m for which x^2+3xy+x+my-m has two linear factors in x and y with integer coefficients
 
2 The length of the sides of a triangle are integrals and it's area is also integer.one side is 21 and the perimeter is 48. Find the smallest side.
 

Grade:12

2 Answers

Vijay Mukati
8 years ago
Dear Student,

For Q1.

From the given two degree equation, we can write that:
x2+ 3xy + x +my -m = (ax+by+c)(dx+e)
Now multiply the two factors on RHS and equate the coefficient on both the sides.

On compairing you will directly get the value of m.

Thanks.
Vijay Mukati
8 years ago
For Q2:

If we consider x to be the shortest side, then other sides will be27-x, 21
Now,x+21 > 27-x or x > 3. (Sum of any two sides is greater then the third side)
And x < 27 - x or x < 14 (Since x is the samllest side)
Therefore, 3 <x<14.

Now apply Heron's Formula.

s = (a+b+c)/2 = 48/2 = 24

A = sqrt(24(24-x)(x-3)(3))
A = sqrt(72(x-3)(24-x))
A = 6sqrt(2(x-3)(24-x))

The possible values of x are from 4 to 13.
Only for x = 10 will give an integeral area.
Therefore the answer is 10.

Thanks